. A corporate note is bought on 5/3/2005 and matures to $4000 on 5/9/2005. On 

5/7/2005 this note is sold to a third party requiring 7%. Find the amount the 

third person paid for the corporate note. Find the actual interest the original 

holder lost on his investment.


1
Expert's answer
2021-05-02T08:48:58-0400

The third person paid:

"Price= Buying price\times1.07"

Current value of note:

Assume it is X, so, grow X for six months to give you $4,000

"X(1.07)^{0.5}=4,000"

"1.22474487X=4000, X=3266"

Price paid by third person:

"Price= 3266\times1.07=3,495"

Third person paid $ 3,495.

Total actual interest as at selling date.

Total interest=4,000-3,266=$734

"Monthly = \frac{734}{6}=122.33\times4=489"


Interest lost= accumulated- on selling

"Interest lost=489-(3495-3266)=260"

Interest lost is $260